The length of a rectangular garden is 3 meters more than its width. If the perimeter of the garden is 50 meters, what is the area of the rectangle? Thankyou. :)
Found 2 solutions by unlockmath, Leaf W.:
Answer by unlockmath(1688)

Hello,
We can set this up as:
(x+3)+ (x+3)+ x + x = 50
Combine like terms:
4x + 6 = 50
Subtract 6 and divide by 4:
x=11
So the width is 11 meters and length is 14 meters.
Now we can find the are to be:
154 Sq meters.
Make sense?
RJ
www.math-unlock.com

Answer by Leaf W.(135)

Hi!
*
Let us call the width of the garden (in meters) "x."
*
Since the length is 3 meters greater than the width, the length = x + 3.
*
Now form an equation using these expressions to show the perimeter of the garden:
Perimeter = 2*width + 2*length
Perimeter = 2x + 2(x + 3)
*
As we know that the perimeter is 50 meters, we can plug this in and solve for x:
50 = 2x + 2(x + 3)
50 = 2x + 2x + 6
44 = 4x
4x = 44
x = 11
*
Therefore, the width = 11 meters. To solve for the length, just plug x = 11 into the expression for length: length = x + 3 = 11 + 3 = 14.
*
Now we know that the width = 11 meters and the length = 14 meters.
*
To find the area, just plug these values into the area formula for a rectangle:
Area = length*width
Area = 14*11
Area = 154
*
Therefore, the area of the rectangular garden is 154 square meters. =)